OpenStudy (anonymous):

Which of the following represents this function written in intercept form? y = -x^2+x+12 A. y = -(x+2)(x+6) B. y = -(x+3)(x-4) C. y = (-x+3)(x-4) D. y = -(x+2)(x-6)

OpenStudy (anonymous):

@KeithAfasCalcLover do you know if this would be A? :)

OpenStudy (anonymous):

bye?

OpenStudy (anonymous):

LOL...that actually was a pleasant surprise comment lol "Ali I leave you alone for couple minutes and look what you have gotten yourself into" lol

OpenStudy (anonymous):

Real mature SwolezNATION

OpenStudy (anonymous):

hahah no kidding! I was like "reallllly" -_-

OpenStudy (anonymous):

Anyways ehh back to the question haha

OpenStudy (anonymous):

I believe the answer is B Guys are we here for swolenation's ehh...libido, or for math?!

OpenStudy (anonymous):

THANK YOU @KeithAfasCalcLover ! :)

OpenStudy (anonymous):

How did you get B as the answer?

OpenStudy (anonymous):

ali1029 it is B I believe

OpenStudy (anonymous):

Are you in 1st grade?

OpenStudy (anonymous):

Whoa. At first, it just brought a smirk to my face, now its really rude swole nation. Somebody, ban him please

OpenStudy (anonymous):

It'll be taken care of. I can assure you.

OpenStudy (anonymous):

Yes please ban, I dont even know this immature person. & thanks Devin & Keith!

OpenStudy (anonymous):

\[-(x+3)(x-4)=-(x^2-4x+3x-12)=-(x^2-x-12)=-x^2+x+12\]

OpenStudy (anonymous):

thisIZnotSWOLENATION! I hope you're using a Proxy or a VPN.

OpenStudy (anonymous):

Oh I see, the negative signs kept confusing me! thank you for explaining :)

OpenStudy (anonymous):

thisIZnotSWOLENATION! Ya know what an IP address is?

OpenStudy (anonymous):

LOL @devinthedev Haha, you really came down hard eh?

OpenStudy (anonymous):

Hahah Devin scared him away! Thank you!

OpenStudy (anonymous):

It just bothers me when people are rude. Ya know?

OpenStudy (anonymous):

Yes I completely get it. I really appreciate your help in getting rid of that jerk!

OpenStudy (anonymous):

Me too. But it does appear like he is spamming Openstudy. I was helping another girl with math and a really similar thing is going on...

OpenStudy (anonymous):

I'm just here to try and learn math lol

OpenStudy (anonymous):

Keith, mind sending me a link?

OpenStudy (anonymous):

Well that's a noble intention @ali1029 haha

OpenStudy (anonymous):

Better yet @devinthedev , check this out: http://openstudy.com/users/swolenationreturnz!!

OpenStudy (anonymous):

Haha seriously! I don't get why people use this to meet people and talk about life and get internet relationships lol :P

OpenStudy (anonymous):

LOL thats like his 3rd account!

OpenStudy (anonymous):

Haha that's what im counting. At least his persistence is admirable haha

OpenStudy (anonymous):

Hahaha! I love how every one of his accounts has swole nation in it..

OpenStudy (anonymous):

I need to be promoted to website security analyst. lol. This site has a ton of vulns.

OpenStudy (anonymous):

Well my vote would definitely be with you!

OpenStudy (anonymous):

So would mine! :) And well I mean if you get a relationship over the internet, good for you, like I don't know what else to say. But if you're all rude about it saying...[insert #swolenation comment] haha then that's a COMPLETE different story. Besides, there's places to talk about life. It's called twitter or facebook...just saying.

OpenStudy (anonymous):

I really could contribute to this community quite well as the security analyst. Even if it was a volunteer position.

OpenStudy (anonymous):

I'm going to look for a Dox

OpenStudy (anonymous):

EXACTLY! Hahaha that's more so how I feel. & You should definitely look into it Dev :)

OpenStudy (anonymous):

So what were all of his user accounts?

OpenStudy (anonymous):

I agree Dev. You have proved your potential already haha

OpenStudy (anonymous):

swolenationreturnz!!, thisIZnotSWOLENATION!, SwolezNATION?

OpenStudy (anonymous):

I wish there were some mods or the admin was online

OpenStudy (anonymous):

I got : @swolenation, @thisIZnotSWOLENATION! and @SwoleNationreturnzz

OpenStudy (anonymous):

SwolezNATION, thisIZnotSWOLENATION! & swolenationreturnz!!

OpenStudy (anonymous):

Who is the admin?

OpenStudy (anonymous):

Not sure, it would help if you guys vouched for me though :)

OpenStudy (anonymous):

Sure, what can I do to help?

OpenStudy (anonymous):

Alright, I have a mod coming over right now

OpenStudy (anonymous):

OH, this should be good. Haha ps. I don't know about you guys, but you know what makes me laugh? When people put things like "PLZ Answer my question, there's a MEDAL in it for ya!!!" LOL Like you think if there's no medal I wont help?

OpenStudy (anonymous):

Hahaha yeah, I just like giving out medals because it's an extended thank you to me :p

OpenStudy (anonymous):

Haha I agree. It's a great feeling!

OpenStudy (anonymous):

ali1029, Sorry about what happened. We'll definitely get it sorted out.

OpenStudy (anonymous):

I wish I could give out multiple medals :(

OpenStudy (anonymous):

Its okay Devin, thank you so much for helping me out, I really appreciate it. :)

OpenStudy (anonymous):

Haha I do too, I think that would be great & Thanks DEV!

OpenStudy (anonymous):

I'll speak with the admin and will try to get that IP addy inputted into some spam databases.

OpenStudy (anonymous):

Thank you so much! You're awesome Dev!

OpenStudy (anonymous):

COOL! Nice man. Im supporting you ALL The way ;)

OpenStudy (anonymous):

*ahem. So ehh this question sounds like it's from grade ten no?

OpenStudy (anonymous):

My math question? hahah I'm a senior but I'm taking pre cal ._.

OpenStudy (anonymous):

The main site will be http://www.stopforumspam.com/ it'll block his use of sites that use the SFS API. Which are a ton nowadays.

OpenStudy (anonymous):

Woohoo!

OpenStudy (anonymous):

Haha cool!

OpenStudy (anonymous):

Pre calc...that's grade 12?

OpenStudy (anonymous):

I guess not, I was supposed to take Geometry my freshman year but I ended up taking Algebra 1 because my schedule messed up so now I'm taking Pre Cal & math is like my kryptonite so that doesn't help at all :p

OpenStudy (anonymous):

I'll be back in 10.

OpenStudy (anonymous):

Sounds good Dev! I'm signing off for a while but if theres anything else you need let me know & I'll get back to you as soon as I can :) Thanks again for everything!

OpenStudy (anonymous):

I'm back, any of you still here?

OpenStudy (anonymous):

Hello Luigi0210

OpenStudy (anonymous):

@AccessDenied @thomaster

OpenStudy (anonymous):

@amistre64 @hartnn